OpenStudy (anonymous):

Give a unique example of each: 1. Part 1: An absolute value equation with two solutions 2. Part 2: An absolute value equation with one solution 3. Part 3: An absolute value equation with no solutions

OpenStudy (anonymous):

\[\left| x \right| = 4\] (or any positive number) \[\left| x \right| = 0\] \[\left| x \right| = -4\] (or any negative number)